OpenStudy (anonymous):

What is the 8th number in the pattern: 20, 12, 4,-4, …?

OpenStudy (anonymous):

It's going down by 8 so if we continue it it would just be -12,-20,-28,-36<== Your answer.

OpenStudy (anonymous):

= a+(n-1)d =20+(8-7)(-8) =20+7(-8) =20-56 =-36

OpenStudy (anonymous):

A.-12 B.-28 C.-36 D.-44 ????????????????????

OpenStudy (anonymous):

ahhh...thats what my friend told me to do

OpenStudy (anonymous):

I had studied it in Arithmetic Progressions.